A downtown employee is looking for the best option for parking a car during a 5 day work week. One parking garage offers unlimited parking at a flat rate of $124 per month. Another garage offers a $4 hourly rate for parking. Which option is best for the employee on a monthly basis if the employee is parking in the garage for a total of 160 hours a month?
Solve the system of equations algebraically to determine where the two cost options are equivalent, showing all work.

Answer by Theo(13342)

flat rate is $124 per month.
hourly rate is $4.00 per hour for 160 hours a month.
$4.00 * 160 = $640.00 per month.
flat rate is way better.
the rate for option a is $124.00
the rate for option b is $4.00 * x.
these options will break even when $4.00 * x = $124.00
this occurs when x = $124.00 / $4.00 = 31.
if the employee only parked 31 hours per month, the options would be the same.
option a would be $124.00
option b would be $4.00 * 31 = $124.00